OpenStudy (anonymous):

cos^2 theta + cos theta = 2 cos theta Thank you!

OpenStudy (anonymous):

\[\cos^2(\theta)+\cos(\theta)=2\cos(\theta)\] Do you want to find theta? First combine the like terms \[\cos^2(\theta)-\cos(\theta)=0\] After this it can be easily solved by factoring

OpenStudy (anonymous):

thats the thing lol.... i dont factor tan sin and cos very well:)